# Globalization of Partial Actions of Ordered Groupoids on Rings

### v1

@@ 4. Uniqueness
In this section we will give a sufficient condition for a P.O. action to have a unique globalization. In order to achieve this goal, we introduce the notion of a minimal globalization.
@@
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a preunital P.O. action of $\mathcal{G}$ on a ring $A$ and $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ be a globalization of $\alpha$ on a ring $B$. We say that $\beta$ is a minimal globalization if
@@
Moreover, we introduce a specific type of P.O. action. Given an ordered groupoid $\mathcal{G}$, we define the pseudoproduct *** as
@@
Definition 4.2.
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a P.O. action of $\mathcal{G}$ on a ring $A$. We say that $\alpha$ is a strong P.O. action [16, Definition 3.7] if
@@
    $\alpha_{g}\circ\alpha_{h}=\alpha_{g*h}\circ\text{Id}_{A_{h^{-1}}}$, for all $g,h\in\mathcal{G}$ such that $\exists g*h$.
@@
Since (PS) holds, notice that $\text{dom}(\alpha_{g}\circ\alpha_{h})=\text{dom}(\alpha_{g*h}\circ\text{Id}_{A_{h^{-1}}})$, for all $g,h\in\mathcal{G}$ such that $\exists g*h$. But this is the same as saying that $\alpha_{g}(A_{g^{-1}}\cap A_{h})=A_{g}\cap A_{g*h}$.
@@
(i): If $e\leq r(g)$, then $\alpha_{e}\circ\alpha_{g}=\alpha_{e*g}\circ\text{Id}_{A_{g^{-1}}}=\alpha_{(e|g)}\circ\text{Id}_{A_{g^{-1}}}=\alpha_{(e|g)}$.
Now $A_{(e|g)}=\text{Im}(\alpha_{(e|g)})=\text{Im}(\alpha_{e}\circ\alpha_{g})=\alpha_{e}(A_{e}\cap A_{g})=A_{e}\cap A_{g}$.
(ii): Notice that $(g|e)=(r(g|e)|g)$ by the uniqueness of restrictions. Hence $A_{(g|e)}=A_{(r(g|e)|g)}=A_{g}\cap A_{r(g|e)}$.
(iii): From (PS) it follows that $\alpha_{e}\circ\alpha_{f}=\alpha_{e*f}\circ\alpha_{f}=\alpha_{e\wedge f}\circ\alpha_{f}=\alpha_{e\wedge f}$. Therefore $A_{e\wedge f}=\text{Im}(\alpha_{e\wedge f})=\text{Im}(\alpha_{e}\circ\alpha_{f})=\alpha_{e}(A_{e}\cap A_{f})=A_{e}\cap A_{f}.$
@@
In the case that $\exists(g*h)*k$ and $\exists g*(h*k)$, we have by [20, Lemma 4.1.6] that these pseudoproducts coincide. Inductive groupoids are always pseudoassociative.
@@
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a strong P.O. action of a pseudoassociative groupoid $\mathcal{G}$ on a ring $A$. Then $\alpha$ admits a minimal globalization if, and only if, $\alpha$ is unital. Furthermore, the minimal globalization is unique up to isomorphisms.
@@
Let $B_{g}$ be the ideal generated by $\bigcup_{r(h)=r(g)}\gamma_{h}(\psi_{d(h)}(A_{d(h)}))$, for all $g\in\mathcal{G}$. Notice that $B_{g}\subseteq F_{g}$. Consider $B=\sum_{e\in\mathcal{G}_{0}}B_{e}$. Define $\beta_{g}:=\gamma_{g}|_{B_{g^{-1}}}$, for all $g\in\mathcal{G}$.
@@
(i): Let $e\in\mathcal{G}_{0}$, $h\in\mathcal{G}$ be such that $r(h)=e$, $a\in A_{d(h)}$ and $b\in A_{e}$. It is enough to prove that $\psi_{e}(b)\beta_{h}(\psi_{d(h)}(a)),\beta_{h}(\psi_{d(h)}(a))\psi_{e}(b)\in\psi_{e}(A_{e})$. For that, consider $k\in\mathcal{G}$. We have that $\beta_{h}(\psi_{d(h)}(a))\psi_{e}(b)|_{k}=\beta_{h}(\psi_{d(h)}(a))|_{k}\psi_{e}(b)|_{k}$. On the one hand,
@@
The uniqueness now follows by [4, Theorem 2.1]. ∎
@@ 5. Applications
We conclude by presenting two applications of the results of the previous sections. We first show that the partial skew ordered groupoid ring is Morita equivalent to the skew ordered groupoid ring relative to its globalization. Afterwards, as a consequence of the ESN Theorem, we provide a necessary and sufficient condition for a partial action of an inverse semigroup to have a (unique) minimal globalization.
@@ 5.1. A Morita Context
To simplify the notation, in this subsection we will assume that $\varphi_{e}=\iota_{e}:A_{e}\to B_{e}$ is the usual inclusion, so that $A_{e}$ is an ideal of $B_{e}$, for all $e\in\mathcal{G}_{0}$.
@@
Let $\alpha=(A_{g},\delta_{g})_{g\in\mathcal{G}}$ be a unital P.O. action of the ordered groupoid $\mathcal{G}$ on a ring $A$ and $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ be a globalization of $\alpha$ on a ring $B$. Let $R=A\ltimes_{\alpha}^{o}\mathcal{G}$ and $T=B\ltimes_{\beta}^{o}\mathcal{G}$. If $\mathcal{G}_{0}$ is finite,
@@
First notice that the crossed products $A\ltimes_{\alpha}\mathcal{G}$ and $B\ltimes_{\beta}\mathcal{G}$ are associative, so $R$ and $T$ are defined. The proof of (i)-(iii) are similar to the nonordered case [4, Proposition 3.1]. In order to prove (iv), it suffices to show that $T\subseteq 1_{R}T1_{R}$. Since for all $h\in\mathcal{G}$ we have that $B_{h}=\sum_{r(g)\leq r(h)}\beta_{g}(A_{d(g)})$, the result follows from
@@
A Morita context is a sextuple $(R,R^{\prime},M,M^{\prime},\varphi,\varphi^{\prime})$, where $R,R^{\prime}$ are unital rings, $M$ is a $(R,R^{\prime})$-bimodule, $M^{\prime}$ is a $(R^{\prime},R)$-bimodule and $\varphi:M\otimes_{R^{\prime}}M^{\prime}\to R$ and $\varphi^{\prime}:M^{\prime}\otimes_{R}M\to R^{\prime}$ are bimodule morphisms such that:
@@
According to [29, Theorems 4.1.4 and 4.1.17], if $\varphi$ and $\varphi^{\prime}$ are surjective, then the categories of left $R$-modules and of left $R^{\prime}$-modules are equivalent and the rings $R$ and $R^{\prime}$ are said to be Morita equivalent. In this case, we also say that the Morita context $(R,R^{\prime},M,M^{\prime},\varphi,\varphi^{\prime})$ is strict.
@@
The rings $R$ and $T$ as above are Morita equivalent.
@@
Consider $M=1_{R}T$ and $N=T1_{R}$. Clearly $M$ is a $(R,T)$-bimodule and $N$ is a $(T,R)$-bimodule. Defining the maps $\varphi:M\otimes_{T}N\to R$ and $\varphi^{\prime}:N\otimes_{R}M\to T$ by $\varphi(m\otimes n)=mn$ and $\varphi^{\prime}(n\otimes m)=nm$, we have that $(R,T,M,N,\varphi,\varphi^{\prime})$ is a strict Morita context by Proposition 5.1. ∎
@@ 5.2. Partial Actions of Inverse Semigroups
Let $S$ be an inverse semigroup. According to [7], a partial action of an inverse semigroup $S$ on a ring $A$ is a collection $\alpha=(A_{s},\alpha_{s})_{s\in S}$ of ideals $A_{ss^{-1}}$ of $A$, ideals $A_{s}$ of $A_{ss^{-1}}$ and ring isomorphisms $\alpha_{s}:A_{s^{-1}}\to A_{s}$ such that
  1. (P1)
    $A=\sum_{e\in E(S)}A_{e}$.
  2. (P2)
    $\alpha_{s}(A_{s^{-1}}\cap A_{t})=A_{s}\cap A_{st}$, for all $s,t\in S$.
  3. (P3)
@@
By [17, Theorem 6.1, Lemma 6.2], we have that the partial actions of an inverse semigroup $S$ on a ring $A$ are in one-to-one correspondence with the strong P.O. actions of an associated inductive groupoid $\mathbb{G}(S)$ on $A$. The ESN Theorem [20, Theorem 4.1.8] also ensures that the global actions of $S$ are in one-to-one correspondence with the global actions of $\mathbb{G}(S)$.
These remarks enable us to translate the theory that we have developed so far to the case of inverse semigroups. More precisely, we have the following result.

### v2

@@ 4. Uniqueness
In this section, we provide a sufficient condition for a P.O. action to have a unique globalization. In order to achieve this goal, we introduce the notion of a minimal globalization.
@@
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a P.O. action of $\mathcal{G}$ on a ring $A$ and $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ be a globalization of $\alpha$ on a ring $B$. We say that $\beta$ is a minimal globalization if
@@
Further on in Example 4.8 we will see some examples of minimal ordered globalizations in contrast with ordered globalizations. Besides that, we need to study some properties of strong partial actions.
Given an ordered groupoid $\mathcal{G}$, we define the pseudoproduct * as
@@
Lemma 4.2.
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a strong P.O. action of $\mathcal{G}$ on a ring $A$. Then it holds
@@
    $\alpha_{g}\circ\alpha_{h}=\alpha_{g*h}\circ\text{Id}_{A_{h^{-1}}}$, for all $g,h\in\mathcal{G}$ such that $g*h$ is defined.
Conversely, if $\alpha$ is a P.O. action that satisfies (PS), then $\alpha$ is strong.
Proof.
Given $(g,h)\in\mathcal{G}_{2}$, we have that
$\alpha_{g}\circ\alpha_{h}$ $=\alpha_{g}\circ\alpha_{h}|_{\alpha_{h^{-1}}(A_{g^{-1}}\cap A_{h})}=\alpha_{g}\circ\alpha_{h}|_{A_{h^{-1}}\cap A_{(gh)^{-1}}}=\alpha_{g}\circ\alpha_{h}\circ\text{Id}_{A_{(gh)^{-1}}\cap A_{h^{-1}}}$
$=\alpha_{gh}\circ\text{Id}_{A_{(gh)^{-1}}\cap A_{h^{-1}}}=\alpha_{gh}\circ\text{Id}_{A_{(gh)^{-1}}}\circ\text{Id}_{A_{h^{-1}}}=\alpha_{gh}\circ\text{Id}_{A_{h^{-1}}}.$
Now, if $g,h\in\mathcal{G}$ are such that the pseudoproduct $g*h$ is defined, that is, such that the meet $e:=d(g)\wedge r(h)$ is defined, then $\alpha_{g}\circ\alpha_{h}=\alpha_{g}|_{A_{h}\cap A_{g^{-1}}}\circ\alpha_{h}|_{\alpha_{h^{-1}}(A_{h}\cap A_{g^{-1}})}$.
Observe that $A_{g^{-1}}\cap A_{h}\subseteq A_{d(g)}\cap A_{r(h)}=A_{e}$, since $\alpha$ is strong. Then
$A_{h}\cap A_{g^{-1}}$ $=A_{h}\cap A_{g^{-1}}\cap A_{e}=A_{h}\cap A_{e}\cap A_{g^{-1}}\cap A_{e}=A_{(e|h)}\cap A_{(e|g^{-1})}=A_{(e|h)}\cap A_{(g|e)^{-1}}.$
Since $\alpha_{(g|e)}=\alpha_{g}|_{A_{(g|e)^{-1}}}$ and $\alpha_{(e|h)}=\alpha_{h}|_{A_{(e|h)^{-1}}}$ we obtain
$\alpha_{g}\circ\alpha_{h}$ $=\alpha_{g}|_{A_{(e|h)}\cap A_{(g|e)^{-1}}}\circ\alpha_{h}|_{\alpha_{(e|h)^{-1}}(A_{(e|h)}\cap A_{(g|e)^{-1}})}$
$=\alpha_{(g|e)}|_{A_{(e|h)}\cap A_{(g|e)^{-1}}}\circ\alpha_{(e|h)}|_{\alpha_{(e|h)^{-1}}(A_{(e|h)}\cap A_{(g|e)^{-1}})}=\alpha_{(g|e)}\circ\alpha_{(e|h)}$
$=\alpha_{(g|e)(e|h)}\circ\text{Id}_{A_{(e|h)^{-1}}}=\alpha_{g*h}\circ\text{Id}_{A_{(e|h)^{-1}}}=\alpha_{g*h}\circ\text{Id}_{A_{(h^{-1}|e)}}.$
By the uniqueness of corestriction we have that $(h^{-1}|e)=(r(h^{-1}|e)|h^{-1})$. In this way, $A_{(h^{-1}|e)}=A_{(r(h^{-1}|e)|h^{-1})}=A_{r(h^{-1}|e)}\cap A_{h^{-1}}$, since $\alpha$ is strong. Now, $r(h^{-1}|e)=r((e|h)^{-1})=d(e|h)$. Hnece $A_{(h^{-1}|e)}=A_{d(e|h)}\cap A_{h^{-1}}$. However $d(g*h)=d(e|h)$, from where $A_{(g*h)^{-1}}\subseteq A_{d(e|h)}$. Therefore
$\alpha_{g}\circ\alpha_{h}$ $=\alpha_{g*h}\circ\text{Id}_{A_{d(e|h)}}\circ\text{Id}_{A_{h^{-1}}}=\alpha_{g*h}\circ\text{Id}_{A_{(g*h)^{-1}}}\circ\text{Id}_{A_{d(e|h)}}\circ\text{Id}_{A_{h^{-1}}}$
$=\alpha_{g*h}\circ\text{Id}_{A_{(g*h)^{-1}}}\circ\text{Id}_{A_{h^{-1}}}=\alpha_{g*h}\circ\text{Id}_{A_{h^{-1}}},$
which is precisely (PS).
For the converse, it only takes us to notice that if $e\leq r(g)$, then
$\alpha_{e}\circ\alpha_{g}$ $=\alpha_{e*g}\circ\text{Id}_{A_{g^{-1}}}=\alpha_{(e|g)}\circ\text{Id}_{A_{g^{-1}}}=\alpha_{(e|g)}.$
Now, $A_{(e|g)}=\text{Im}(\alpha_{(e|g)})=\text{Im}(\alpha_{e}\circ\alpha_{g})=\alpha_{e}(A_{e}\cap A_{g})=A_{e}\cap A_{g}$, showing that $\alpha$ is strong. ∎
@@
When $\alpha$ is strong, notice that $\text{dom}(\alpha_{g}\circ\alpha_{h})=\text{dom}(\alpha_{g*h}\circ\text{Id}_{A_{h^{-1}}})$, for all $g,h\in\mathcal{G}$ such that $\exists g*h$. But this is the same as saying that $\alpha_{g}(A_{g^{-1}}\cap A_{h})=A_{g}\cap A_{g*h}$.
@@
(i): Notice that $(g|e)=(r(g|e)|g)$ by the uniqueness of restrictions. Hence $A_{(g|e)}=A_{(r(g|e)|g)}=A_{g}\cap A_{r(g|e)}$.
(ii): From (PS) it follows that $\alpha_{e}\circ\alpha_{f}=\alpha_{e*f}\circ\alpha_{f}=\alpha_{e\wedge f}\circ\alpha_{f}=\alpha_{e\wedge f}$. Therefore $A_{e\wedge f}=\text{Im}(\alpha_{e\wedge f})=\text{Im}(\alpha_{e}\circ\alpha_{f})=\alpha_{e}(A_{e}\cap A_{f})=A_{e}\cap A_{f}.$
@@
In the case that $\exists(g*h)*k$ and $\exists g*(h*k)$, we have by [26, Lemma 4.1.6] that these pseudoproducts coincide. Inductive groupoids are always pseudoassociative.
Now we have all the tools needed to characterize the uniqueness of minimal ordered globalizations.
@@
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a preunital, strong P.O. action of a pseudoassociative groupoid $\mathcal{G}$ on a ring $A$. Then $\alpha$ admits a minimal globalization if and only if $\alpha$ is unital. Furthermore, the minimal globalization is unique up to equivalence.
@@
Let $B_{g}$ be the subring of $F_{g}$ generated by $\bigcup_{r(h)=r(g)}\gamma_{h}(\psi_{d(h)}(A_{d(h)}))$, for all $g\in\mathcal{G}$. Notice that $B_{g}\subseteq F_{g}$ and that $B_{g}$ may not be unital. Consider $B=\sum_{e\in\mathcal{G}_{0}}B_{e}$. Define $\beta_{g}:=\gamma_{g}|_{B_{g^{-1}}}$, for all $g\in\mathcal{G}$.
@@
(i): Let $e\in\mathcal{G}_{0}$, $h\in\mathcal{G}$ be such that $r(h)=e$, $a\in A_{d(h)}$ and $b\in A_{e}$. It is enough to prove that $\psi_{e}(b)\beta_{h}(\psi_{d(h)}(a)),\beta_{h}(\psi_{d(h)}(a))\psi_{e}(b)\in\psi_{e}(A_{e})$. For that, take $k\in\mathcal{G}$. We have that $\beta_{h}(\psi_{d(h)}(a))\psi_{e}(b)|_{k}=\beta_{h}(\psi_{d(h)}(a))|_{k}\psi_{e}(b)|_{k}$. On the one hand,
@@
The uniqueness now follows by [5, Theorem 2.1]. ∎
@@ 5. Applications
We conclude by presenting two applications of the results of the previous sections. We first show that the partial skew ordered groupoid ring is Morita equivalent to the skew ordered groupoid ring relative to its globalization. Afterwards, as a consequence of the ESN Theorem [26, Theorem 4.1.8], we provide a necessary and sufficient condition for a partial action of an inverse semigroup to have a (unique) minimal globalization.
@@ 5.1. A Morita Context
To simplify the notation, in this subsection we will assume that $\varphi_{e}=\iota_{e}:A_{e}\to B_{e}$ is the usual inclusion, so that $A_{e}$ is an ideal of $B_{e}$, for all $e\in\mathcal{G}_{0}$. Also, assume that $\mathcal{G}_{0}$ is finite.
@@
Let $\alpha=(A_{g},\delta_{g})_{g\in\mathcal{G}}$ be a unital P.O. action of the ordered groupoid $\mathcal{G}$ on a ring $A$ and $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ be a globalization of $\alpha$ on a ring $B$. Let $R=A\ltimes_{\alpha}^{o}\mathcal{G}$ and $T=B\ltimes_{\beta}^{o}\mathcal{G}$. Then:
@@
First, notice that the crossed products $A\ltimes_{\alpha}\mathcal{G}$ and $B\ltimes_{\beta}\mathcal{G}$ are associative, so $R$ and $T$ are defined. Also, since $\alpha$ is unital and $\mathcal{G}_{0}$ is finite, we have that $1_{R}$ is defined. The proof of (i)-(iii) are similar to the nonordered case [5, Proposition 3.1]. In order to prove (iv), it suffices to show that $T\subseteq 1_{R}T1_{R}$. Since for all $h\in\mathcal{G}$ we have that $B_{h}=\sum_{r(g)\leq r(h)}\beta_{g}(A_{d(g)})$, the result follows from
@@
[…]
According to [19], a Morita context is a sextuple $(R,R^{\prime},M,M^{\prime},\varphi,\varphi^{\prime})$, where $R,R^{\prime}$ are idempotent rings, $M$ is a $(R,R^{\prime})$-bimodule, $M^{\prime}$ is a $(R^{\prime},R)$-bimodule and $\varphi:M\otimes_{R^{\prime}}M^{\prime}\to R$ and $\varphi^{\prime}:M^{\prime}\otimes_{R}M\to R^{\prime}$ are bimodule morphisms such that:
@@
By [19, Proposition 2.6], if ${}_{R}M$, $M^{\prime}_{R}$, $M_{R^{\prime}}$ and ${}_{R^{\prime}}M^{\prime}$ are unital modules and $\varphi$ and $\varphi^{\prime}$ are surjective, then the categories $R$-mod and $R^{\prime}$-mod (resp. mod-$R$ and mod-$R^{\prime}$) are equivalent and the rings $R$ and $R^{\prime}$ are said to be Morita equivalent.
@@
The rings $R=A\ltimes_{\alpha}^{o}\mathcal{G}$ and $T=B\ltimes_{\beta}^{o}\mathcal{G}$ are Morita equivalent.
@@
First note that $R$ is unital, since $\alpha$ is unital and $\mathcal{G}_{0}$ is finite. Hence, $R$ is idempotent. Since $B$ has local units and $\mathcal{G}_{0}$ is finite, we also have that $T$ has local units and therefore is idempotent.
Consider $M=1_{R}T$ and $N=T1_{R}$. Clearly $M$ is a $(R,T)$-bimodule and $N$ is a $(T,R)$-bimodule. Defining the maps $\varphi:M\otimes_{T}N\to R$ and $\varphi^{\prime}:N\otimes_{R}M\to T$ by $\varphi(m\otimes n)=mn$ and $\varphi^{\prime}(n\otimes m)=nm$, we have that this maps are surjective and $(R,T,M,N,\varphi,\varphi^{\prime})$ is a Morita context by Proposition 5.1. It is also easy to see that $M_{T},$ ${{}_{R}M},N_{R},$ and ${{}_{T}N}$ are unital modules since $R$ and $T$ are idempotent rings. ∎
@@ 5.2. Partial Actions of Inverse Semigroups
We will begin this subsection by reminding the reader of the notion of premorphism.
Definition 5.3.
[20, p. 184] Let $\mathcal{G}$ and $\mathcal{H}$ inductive groupoids. A map $\psi:\mathcal{G}\to\mathcal{H}$ is said to be an inductive groupoid premorphism if:
  1. (i)
    $\psi(g)*\psi(h)\leq\psi(gh)$, for all $(g,h)\in\mathcal{G}_{2}$;
  2. (ii)
    $\psi(g)^{-1}=\psi(g^{-1})$, for all $g\in\mathcal{G}$;
  3. (iii)
    $\psi$ preserves orders.
Remark 5.4.
  • 1.
    By [21, Definition 2.6], every inductive groupoid premorphism satisfies $\psi(g)*\psi(h)=\psi(gh)*d(\psi(h))$, for all $(g,h)\in\mathcal{G}_{2}$.
  • 2.
    By [20, Lemma 4.2(b)], we always have $d(\psi(g))\leq\psi(d(g))$, by all $g\in\mathcal{G}$.
  • 3.
    By [21, Definition 2.6], in all inductive groupoid premorphisms we have that if $e\in\mathcal{G}_{0}$ and $g\in\mathcal{G}$ are such that $e\leq d(g)$, then $d(\psi(g|e))=\psi(e)\wedge d(\psi(g))$; and if $e\in\mathcal{G}_{0}$ and $g\in\mathcal{G}$ are such that $e\leq r(g)$, then $r(\psi(e|g))=\psi(e)\wedge r(\psi(g))$.
Every inverse semigroup $S$ with set of idempotents $E(S)$ has a partial natural order $\preceq$ given by:
$s\preceq_{S}t$ if and only if there exists $e\in E(S)$ such that $s=te$.
Definition 5.5.
[22, p. 19] Let $S$ and $T$ be inverse semigroups. A map $\psi:S\to T$ is said to be an inverse semigroup premorphism if, for all $s,t\in S$, it holds:
  1. (i)
    $\psi(s)\psi(t)\preceq\psi(st)$;
  2. (ii)
    $\psi(s)^{-1}=\psi(s^{-1})$;
  3. (iii)
    If $s\preceq t$, then $\psi(s)\preceq\psi(t)$.
The Ehresmann-Schein-Nambooripad Theorem establishes a relation between the categories of inductive groupoids and inverse semigroups. From now on, we will refer to it as the ESN Theorem, written below.
Theorem 5.6 (ESN).
  1. (i)
    [26, Theorem 4.1.8] The category of inductive groupoids and inductive functors and the category of inverse semigroups and homomorphisms are isomorphic.
  2. (ii)
    [22, Theorem 6.1] The category of inductive groupoids and inductive groupoid premorphisms and the category of inverse semigroups and inverse semigroup premorphisms are isomorphic.
Let $S$ be an inverse semigroup, and $E(S)$ the subset of idempotent elements of $S$. We adopt the definition of partial inverse semigroup action as stated in [9, Definition 2.1] with a slight weakening of the original assumptions. Here, we assume that $A_{s}$ is an ideal of $A_{ss^{-1}}$ instead of asking that $A_{s}$ is an ideal of $A$, for each $s\in S$.
Definition 5.7.
A partial action of an inverse semigroup $S$ on a ring $A$ is a collection $\alpha=(A_{s},\alpha_{s})_{s\in S}$ of ideals $A_{ss^{-1}}$ of $A$, ideals $A_{s}$ of $A_{ss^{-1}}$ and ring isomorphisms $\alpha_{s}:A_{s^{-1}}\to A_{s}$ such that
  1. (P1’)
    $A=\sum_{e\in E(S)}A_{e}$;
  2. (P2’)
    $\alpha_{s}(A_{s^{-1}}\cap A_{t})=A_{s}\cap A_{st}$, for all $s,t\in S$;
  3. (P3’)
@@
The definitions of unital and preunital partial actions of inverse semigroups are analogous to the case of ordered groupoids. Next, we define the concept of globalization of a partial inverse semigroup action.
Definition 5.8.
Let $\alpha=(A_{s},\alpha_{s})_{s\in S}$ be a partial inverse semigroup action of an inverse semigroup $S$ on a ring $A$. A global action $\beta=(B_{s},\beta_{s})_{s\in S}$ of $S$ on a ring $B$ is said to be a globalization of $\alpha$ if, for all $e\in E(S)$, there exists a ring monomorphism $\varphi_{e}:A_{e}\to B_{e}$ such that:
  1. (i)
    $\varphi_{e}(A_{e})$ is an ideal of $B_{e}$;
  2. (ii)
    $\varphi_{ss^{-1}}(A_{s})=\varphi_{ss^{-1}}(A_{ss^{-1}})\cap\beta_{s}(\varphi_{s^{-1}s}(A_{s^{-1}s}))$;
  3. (iii)
    $\beta_{s}\circ\varphi_{s^{-1}s}(a)=\varphi_{ss^{-1}}\circ\alpha_{s}(a)$, for all $a\in A_{s^{-1}}$;
  4. (iv)
    $B_{s}=\sum_{tt^{-1}=ss^{-1}}\beta_{t}\left(\varphi_{t^{-1}t}(A_{t^{-1}t})\right)$.
We will denote by $\mathsf{Set}$ the category of sets and functions and by $\mathsf{Ring}$ the category of rings and ring homomorphisms. Given $A,B$ in $\text{Obj}(\mathsf{Ring})$ and $\varphi:A\to B$ in $\text{Mor}(\mathsf{Ring})$, we consider the forgetful functor $F:\mathsf{Ring}\to\mathsf{Set}$, where the image of the ring $A$ is the set $A$ (without the ring structure), and the image of the ring homomorphism $\varphi:A\to B$ is the function of sets $\varphi:A\to B$ (without the homomorphisms properties).
Let $A$ be a ring. We define $\mathcal{I}(A):=\mathcal{I}(F(A))=\{f:I\to J:I,J\subseteq A,f\text{ is a bijection}\}$. With the usual composition, we have that $\mathcal{I}(A)$ is an inverse semigroup.
Let $S$ be an inverse semigroup and consider a partial action $\alpha=(A_{s},\alpha_{s})_{s\in S}$ of $S$ on a ring $A$.
Lemma 5.9.
The partial action $\alpha$ gives rise to a premorphism $\gamma_{\alpha}:S\to\mathcal{I}(A)$, where $\gamma_{\alpha}(s)$ is a bijection between the subsets $\text{dom}(\gamma_{\alpha}(s))=F(A_{s^{-1}})$ and $\text{Im}(\gamma_{\alpha}(s))=F(A_{s})$ of $F(A)$, defined as $\gamma_{\alpha}(s)(a)=\alpha_{s}(a)$, for all $a\in F(A_{s^{-1}})$.
Proof.
Condition (i) of premorphism follows from (P2’) and (P3’) of the definition of partial action. Condition (ii) of the premorphism follows from [9, Proposition 2.2(c)]. Condition (iii) of the premorphism follows from (PO). ∎